Research has shown that approximately 1 woman in

500500

carries a mutation of a particular gene. About

6969 %

of women with this mutation develop

breastbreast

cancer. Find the probability that a randomly selected woman will carry the mutation of this gene and will develop

breastbreast

cancer.
Answer #1

P(carries a mutation of a particular gene) = 1/500 = 0.002

P(breast cancer| carry mutation of a gene) = 0.69

P(woman will carry the mutation of this gene and will develop breast cancer ) = 0.69*0.002=0.00138 (answer)